cubed root of x+2=2
Answer by checkley77(12844)

cubed root of x+2=2
CUBE BOTH SIDES.
X+2=2^3
X+2=8
X=8-2
X=6 ANS.
PROOF:
CUBERT(6X+2)=2
CUBERT(8)=2
2=2